# Angles Don’t Lie: Unlocking Training-Efficient RL Through the Model’s Own Signals

### v1

@@ Abstract
Current Reinforcement Fine-tuning (RFT) paradigms for Large Language Models (LLMs) suffer from sample inefficiency due to the redundant exposure of identical queries under uniform data sampling. While previous work has explored curriculum learning via heuristic difficulty metrics, these strategies exhibit limitations by neglecting the intrinsic learning signals generated by the model itself, thus leading to suboptimal training regimes. In this paper, we identify a model-inherent signal termed angle concentration that effectively reflects an LLM’s capacity to learn from specific data. We theoretically and empirically demonstrate a correlation between the angular distribution of token hidden state vectors and the resulting gradient, revealing a learning preference for data exhibiting higher angle concentration. Inspired by this finding, we propose GAIN-RL, a Gradient-driven Angle-Informed Navigated RL framework. By leveraging the model’s intrinsic angle concentration signal, GAIN-RL dynamically selects training data in each epoch, ensuring consistently impactful gradient updates and thus significantly enhancing overall training efficiency. Empirical evaluations show that GAIN-RL (GRPO) achieves over a 2.5$\times$ acceleration in training efficiency across diverse mathematical and coding tasks and varying model scales. Furthermore, GAIN-RL (GRPO)’s efficient sampling yields data-efficient training, achieving better performance with half the original data compared to vanilla GRPO with full training data. Code is realsed at https://github.com/wangqinsi1/GAINRL/tree/main.
@@ 1 Introdction
We aim to attack the aforementioned challenges and propose a model-informed signal that (1) reflects the learning capacity of a specific model on particular data, (2) incurs minimal computational costs, and (3) maintains generalizability across diverse models and datasets. Achieving such requirements is inherently challenging, as accurately capturing model-data interactions typically needs a resource-intensive decoding stage. To overcome this challenge, we explore three key questions in this work:
@@
Building upon these insights, we propose GAIN-RL, a Gradient-driven Angle-Informed Navigated Reinforcement Learning Framework, illustrated in Fig. 1. GAIN-RL comprises three primary components: Data Reordering, Data Sampling and Probability Update. Before training, we reorder the training data based on model-informed angular concentration to enhance learning efficiency. During training, a dynamic Gaussian probability sampling strategy progressively guides the model towards data with lower angular concentration, with the pace adjusted according to real-time accuracy and angular signals. Notably, our data preprocessing involves only the inexpensive pre-filling stage, requiring under 10 minutes for over 7,000 samples. Overall, GAIN-RL provides a plug-and-play, broadly applicable, and data-efficient reinforcement learning solution. Experiments validate that GAIN-RL accelerates training by over 2.5$\times$ and, remarkably, surpasses full-sample performance using only half the data—demonstrating its powerful capability in driving data-efficient RFT.
@@
Figure 2: Visualization of models’ responses. We evaluate the first 100 GSM8K questions by having Qwen2.5-0.5B-Instruct and LLaMA3.2-1B-Instruct each generate 10 answers per question, recording the number of correct responses. ChatGPT-4o rates question ease on a 0–10 scale (10 = easiest).
2 Model-Informed Data Evaluation Signals
@@ 2.1 Which Signals Should We Focus On?
$\!\!\!\lVert\nabla_{\bm{W}}\mathcal{L}\rVert_{F}^{2}\!=\!\sum_{i=1}^{m}\sum_{j=1}^{m}\bigl{(}\bm{x}_{i}\bm{x}_{j}^{\!\top}\bigr{)}\!\bigl{(}(\nabla_{\bm{a}}\mathcal{L})_{i}(\nabla_{\bm{a}}\mathcal{L})_{j}^{\!\top}\bigr{)}\!=\!\sum_{i=1}^{m}\sum_{j=1}^{m}\|\bm{x}_{i}\|\|\bm{x}_{j}\|\cos\theta_{i,j}\bigl{(}(\nabla_{\bm{a}}\mathcal{L})_{i}(\nabla_{\bm{a}}\mathcal{L})_{j}^{\!\top}\bigr{)},$ (4)
@@ 2.2 What Are the Characteristics of This Signal?
Therefore, assume the length of the input tokens is $m$. The first $n$ tokens constitute the system prompt and the few-shot examples, which are the same across all data samples. The remaining $m-n$ tokens represent the specific question to be answered. The characteristics we should focus on are:
@@
where $\mathcal{C}_{\mathrm{intra}}$ measures angle concentration within the question and $\mathcal{C}_{\mathrm{inter}}$ measures angle concentration between question tokens and the system prompt and few-shot tokens. We measure both at the final layer, where inter-segment clustering is maximal. Concentration within the system prompt and few-shot tokens is omitted, as it is constant across different questions.
@@
Figure 3: Visualization of Layer-wise Concentration Pattern. Experiment is performed on the Qwen2.5-0.5b-Instruct. In each subplot, the pixel at row i, column j represents the cosine similarity of the angle between the i-th and j-th token hidden states of the layer output. Blue, yellow, and gray arrows above the figure represent the tokens of the system prompt, few-shot examples and question, respectively. To better highlight the pattern, values are clipped between the 3rd and 97th percentiles.
Figure 4: Visualization of Epoch-wise Concentration Pattern. (Left) $\mathcal{C}_{\mathrm{intra}}$; (Mid) $\mathcal{C}_{\mathrm{inter}}$; (Right) $\mathcal{C}_{\mathrm{intra}}+\mathcal{C}_{\mathrm{inter}}$. We train Qwen2.5-0.5B-Instruct and LLaMA3.2-1b-Instruct on GSM8K using GRPO for 250 epochs. To accelerate observation, we use a training batch size of 16, generate 4 responses per question, and set the learning rate to 1e-5. After each epoch, we perform pre-filling on the model on the entire dataset and record the angle concentration from the hidden states of the final layer output.
@@
A2. The angles between token hidden states show both intra-segment and inter-segment concentrations during inference. In particular, we should pay attention to the final layer as the inter-segment clustering is most pronounced, resulting in the highest overall concentration.
@@ 2.3 How Can This Signal Be Leveraged to Accelerate Training?
Data-wise Observation. Furthermore, to examine data-wise angle behavior during training, we track the model’s responses to samples with varying angle concentrations over epochs. As depicted in Fig. 5, surprisingly, the results revealed that during training, the model tends to prioritize learning from higher-angle concentration data before addressing lower-angle concentration data, which we introduce as Data-wise Angle Concentration Pattern. For instance, by epoch 100, questions with maximal angular measurements were almost entirely answered correctly, whereas those with smaller angles remained uncorrected. Note that the angle concentration distribution is measured on the untrained model, indicating that despite its evolution during training, the initial angle concentration of the data provides meaningful guidance for the training process.
@@
Gradient-based Explanation. As shown in Eq. 4, gradients are influenced by both angle concentration and loss. Early in training, when losses are relatively uniform across samples, those with higher angle concentration receive stronger gradients and are learned faster. As training continues, angle concentration rises overall: high-angle samples, already mastered, see lower losses; while low-angle samples gain concentration, inherit larger gradients, and are learned next. This creates a natural, angle-driven learning progression. Unlike traditional curriculum learning based on task difficulty, angle concentration offers a more intuitive, model-centric training signal.
@@ 3 GAIN-RL Framework
Probability Update Based on Accuracy and Angular Signals. The Gaussian mean $\mu_{t}$ sets the peak-sampling region. We initialize $\mu_{0}=0$ to prioritize high-angle concentration data, then gradually increase $\mu_{t}$ as the model masters these samples, shifting focus to lower-angle concentration ones. At each step, $\mu_{t}$ is updated from the batch $d^{(t)}$ using its average accuracy and angle concentration,
@@
GAIN-RL(GRPO) 88.09 67.20 43.37 13.33 27.26 16.54 42.63 80 2.50$\times$
LLaMA 3.2 3B Instruct GRPO 74.60 40.20 19.28 6.67 11.70 8.46 26.8 200 1$\times$
@@
GAIN-RL(GRPO) 76.04 42.00 21.69 6.67 14.22 10.29 28.5 80 2.50$\times$
Qwen 2.5 Math 7B Instruct GRPO 91.96 68.40 40.96 10.00 25.78 20.22 42.89 200 1$\times$
@@
GAIN-RL(GRPO) 93.71 72.80 45.78 13.33 26.81 23.53 46.33 70 2.86$\times$
@@
Qwen 2.5 Coder 3B Instruct GRPO 10.8 21.5 5.15 17.8 78.3 26.7 200 1$\times$
GAIN-RL(GRPO) 12.8 25.1 6.14 18.2 81.5 28.8 110 1.81$\times$
Appendix F.1 experiments reveal that weighting the signal as $\mathcal{C}_{\mathrm{intra}}^{M}+k\cdot\mathcal{C}_{\mathrm{inter}}^{M}$ can further boost performance. For simplicity and broad applicability, we adopt the unweighted signal here and leave signal optimization to future work. Our main goal is to highlight angle concentration as a key signal.
Figure 7: Learning Dynamics of Different Methods on (Left) Qwen2.5-Math-1.5b-Instruct and (Right) LLaMA-3.2-3b-Instruct. The y-axis shows average accuracy over GSM8K, Math, AMC 23, AIME 24, OlympiadBench, and Minerva Math. Performance is evaluated every 5 epochs.
@@
Figure 9: Fine-tuning Performance on Single Tasks. (Left) GSM8k. (Mid) Math. (Right) AMC 23. Models are trained on the training sets and evaluated on their validation sets. ADARFT is excluded from GSM8K due to missing difficulty coefficients.
@@ 4.1 Training Efficiency of GAIN-RL
Hardware Efficiency. To evaluate hardware efficiency, we report the number of epochs required for each method to reach the performance of vanilla GRPO at 200 epochs. As shown in Tab. 1 and Tab. 2, GAIN-RL(GRPO) requires approximately only half the epochs across various models and tasks. Specifically, for the Qwen2.5-Math-1.5B-Instruct and LLaMA3.2-3B-Instruct models, GAIN-RL(GRPO) achieves approximately 2.5$\times$ speedup, significantly outperforming ADARFT(GRPO) speedups of 1.33$\times$ and 1.43$\times$, respectively. Fig. 7 further visualizes performance trajectories during training, where GAIN-RL(GRPO) consistently demonstrates faster convergence and superior performance at every epoch. As analyzed in Sect. 2, the acceleration mainly stems from the ability of our method to maintain strong gradient signals throughout training, leading to faster learning.
Figure 10: (Left) Ablation Study of Module Components. (Right) Small Batch Scalability Test. Experiments use Qwen2.5-0.5b-Instruct on Math training set, evaluate on the test set every 20 epochs.
@@ 4.2 Data Efficiency of GAIN-RL
Table 3: Performance of GAIN-RL combined with PPO. Qwen2.5-0.5b-Instruct is trained and evaluated on three datasets. Implementation details are in Appendix E.
@@ 4.3 RL Algorithms Generalization
To verify the algorithm generalization of GAIN-RL, we also combine GAIN-RL with PPO. We evaluate the final performance and hardware efficiency. Results in Tab. 4.2 highlight substantial gains in both performance and hardware efficiency with GAIN-RL(PPO), achieving an average of 2.2$\times$ training speedup across three mathematical benchmarks. This consistent benefit arises because gradient updates are fundamental across RL algorithms, validating GAIN-RL’s universal acceleration capability.
@@ 4.4 Performance on Individual Tasks
To further demonstrate the generality of GAIN-RL(GRPO), we also evaluate its performance on single-task RFT, which demands higher precision in data ordering and selection due to narrower difficulty ranges. Using Qwen2.5-0.5b-Instruct, we fine-tuned separately on GSM8K, MATH, and AMC training sets and evaluated on their test sets. Results in Fig. 9 indicate that GAIN-RL(GRPO) consistently yielded higher performance and efficiency. Specifically, on GSM8K, GAIN-RL(GRPO) achieved a 3.33$\times$ training speedup and a 4.72% final accuracy improvement. On the more challenging MATH and AMC 23 datasets, GAIN-RL(GRPO) also achieves 2.5$\times$ and 2$\times$ speedups, respectively. In contrast, ADARFT(GRPO) provided less improvement due to its fixed difficulty scoring, which may not align precisely with actual model-perceived difficulty. In contrast, by leveraging model-informed angle signals, GAIN-RL(GRPO) can predict learning priorities more accurately.
@@ 4.5 Ablation Studies
Small-Batch Scalability Test. To validate the scalability of GAIN-RL under reduced training batch sizes, we evaluated the model’s performance with varying batch sizes. Specifically, we conducted experiments using the Qwen2.5-0.5b-instruct model on the Math dataset with training batch sizes of 512, 768, and 1024, respectively. The experimental outcomes, depicted in Fig. 10 (right), illustrate consistent and stable performance improvement across epochs for all batch sizes. Remarkably, even with the batch size reduced to half of the original size (n=512), which significantly reduces computational time and memory usage per epoch, GAIN-RL maintained performance comparable to vanilla GRPO. These results demonstrate that GAIN-RL effectively scales with smaller batch sizes, offering the flexibility to accelerate training by lowering batch size, with only minor performance degradation, especially beneficial under limited computational or memory resources.
@@ 5 Conclusion
We propose GAIN-RL, a novel Reinforcement Learning Fine-tuning framework that leverages the angle concentration signals to dynamically allocate training data at each epoch. GAIN-RL leverage the model’s intrinsic angle concentration signal to dynamically selects training data in each epoch, ensuring consistently impactful gradient updates and thus significantly enhancing overall training efficiency. Furthermore, our empirical results further show that GAIN-RL (GRPO) achieves over a 2.5$\times$ acceleration in training efficiency across diverse mathematical and coding tasks and varying model scales. Overall, GAIN-RL introduces a novel paradigm for training-efficiency RFT, highlighting how model-centric data-processing approaches can remedy the sub-optimality of current RFT methods and further elevate their effectiveness.
@@ Appendix A Related Work
In parallel, data-centric strategies [28, 29, 30, 31, 32, 33] have emerged as promising alternatives for efficient fine-tuning. For example, [28] provides the first quantitative audit of preference datasets, introducing metrics for scale, noise, and information density that expose quality bottlenecks before any policy update. Direct Preference Optimization (DPO) [34] simplifies the entire loop by replacing on-policy RL with a closed-form classification loss, making the choice of high-value preference data the primary driver of alignment quality. To further reduce annotation cost, Active Preference Optimization [35] casts RLHF as an active-learning bandit, adaptively querying only the prompts expected to maximize reward-model improvement. The latest data-centric strategies can be categorized based on their approach to data manipulation: data selection and data sequencing. Data selection techniques involve filtering extensive datasets to retain only a small subset of high-quality training data based on predefined metrics. Methods such as LIMO[12] and s1[13] have demonstrated that carefully curated small supervised fine-tuning datasets can achieve robust performance using orders of magnitude less data. On the other hand, data sequencing strategies[36] enhance model learning speed by rearranging the order of training data within existing datasets. Approaches like ADARFT[14] have shown that dynamically selecting data for each epoch can effectively accelerate the training process.
@@ B.1 Theoretical Justification of Angle-Dependent Effects in LLMs
Figure 11: Visualization of $W_{Q}@W_{K}.T$ at different layers in LLaVA-1.5-7b.
Figure 12: Visualization of $W_{D}@W_{D}.T$ at different layers in LLaVA-1.5-7b.
Figure 13: Visualization of $W_{V}@W_{V}.T$ at different layers in LLaVA-1.5-7b.
@@
For Observation 1, we show $W_{Q}@W_{K}.T$, $W_{V}@W_{V}.T$, $W_{D}@W_{D}.T$ of all different layers of LLaVA-1.5-7B in Fig. 11, 12, and 13. It can be seen that different layers have this orthogonal relationship. In fact, the orthogonal relationship of matrices in neural networks has been studied since a long time ago. In particular, [37] proposed a new regularization method that encourages the weight matrix of the neural network to maintain orthogonality during training by introducing a self-orthogonality module. This method helps to improve the training stability and generalization ability of the model. [38] explores adding orthogonal regularization to weights during training to improve training stability. The author proposed an orthogonal regularization method for weights, aiming to solve the gradient vanishing and explosion problems encountered by deep convolutional neural networks during training. It can be seen that modules with orthogonality are found in various different models to improve the training stability and performance of the model. To the best of our knowledge, we are the first work to intuitively show this orthogonal performance in LLM, which can be more fully explored in subsequent research.
@@
Figure 14: Visualization of $cos(\angle(A_{i},A_{M}))$ and co-act neurons number at different layers in LLaVA-1.5-7b.
@@
Justification. For a single token, suppose its input to the Attention block is $y$. Consider a sequence of tokens $[y_{1},y_{2},\ldots,y_{M}]$, for token $y_{m}$ , its computation in the Attention layer can be expressed as:
@@
$\begin{split}&\|\alpha_{im}V_{i}\|\cos(\angle(V_{i},V_{m}))\\ &\propto(\hat{y}_{i}\,W_{q})\,\bigl{(}\hat{y}_{m}\,W_{k}\bigr{)}^{T}\|V_{i}\|\cos(\angle(V_{i},V_{m}))\\ &=\langle\hat{y}_{i}W_{q},\hat{y}_{m}W_{k}\rangle\langle V_{i},V_{m}\rangle/\|V_{m}\|\\ &=\bigl{(}\hat{y}_{i}(W_{q}W_{k}^{T})\hat{y}_{m}^{T}\bigr{)}\bigl{(}\hat{y}_{i}(W_{v}W_{v}^{T})\hat{y}_{M}^{T}\bigr{)}/\|V_{m}\|\\ \end{split}$ (14)

### v2

# Angles Don’t Lie: Unlocking Training‑Efficient RL Through the Model’s Own Signals

@@ Abstract
[…]
@@ 1 Introdction
We aim to attack the aforementioned challenges and propose a model‑informed signal that (1) reflects the learning capacity of a specific model on particular data, (2) incurs minimal computational costs, and (3) maintains generalizability across diverse models and datasets. Achieving such requirements is inherently challenging, as accurately capturing model-data interactions typically needs a resource-intensive decoding stage. To overcome this challenge, we explore three key questions in this work:
@@
[…]
@@
Figure 2: Visualization of models’ responses. We evaluate the first 100 GSM8K questions by having Qwen2.5‑0.5B‑Instruct and LLaMA3.2‑1B‑Instruct each generate 10 answers per question, recording the number of correct responses. ChatGPT‑4o rates question ease on a 0–10 scale (10 = easiest).
2 Model‑Informed Data Evaluation Signals
@@ 2.1 Which Signals Should We Focus On?
$\lVert\nabla_{\bm{W}}\mathcal{L}\rVert_{F}^{2}\!=\!\sum_{i=1}^{m}\sum_{j=1}^{m}\bigl(\bm{x}_{i}^{\!\top}\bm{x}_{j}\bigr)\!\bigl((\nabla_{\bm{a}}\mathcal{L})_{i}(\nabla_{\bm{a}}\mathcal{L})_{j}^{\!\top}\bigr)\!=\!\sum_{i=1}^{m}\sum_{j=1}^{m}\|\bm{x}_{i}\|\|\bm{x}_{j}\|\cos\theta_{i,j}\bigl((\nabla_{\bm{a}}\mathcal{L})_{i}(\nabla_{\bm{a}}\mathcal{L})_{j}^{\!\top}\bigr),$ (4)
@@ 2.2 What Are the Characteristics of This Signal?
Therefore, assume the length of the input tokens is $m$. The first $n$ tokens constitute the system prompt and the few-shot examples, which are the same across all data samples. The remaining mnm-n tokens represent the specific question to be answered. The characteristics we should focus on are:
@@
where $\mathcal{C}_{\mathrm{intra}}$ measures angle concentration within the question and $\mathcal{C}_{\mathrm{inter}}$ measures angle concentration between question tokens and the system prompt and few-shot tokens. We measure both at the final layer, where inter‑segment clustering is maximal. Concentration within the system prompt and few‑shot tokens is omitted, as it is constant across different questions.
@@
Figure 3: Visualization of Layer-wise Concentration Pattern. Experiment is performed on the Qwen2.5-0.5b-Instruct. In each subplot, the pixel at row i, column j represents the cosine similarity of the angle between the i‑th and j‑th token hidden states of the layer output. Blue, yellow, and gray arrows above the figure represent the tokens of the system prompt, few-shot examples and question, respectively. To better highlight the pattern, values are clipped between the 3rd and 97th percentiles.
Figure 4: Visualization of Epoch-wise Concentration Pattern. (Left) $\mathcal{C}_{\mathrm{intra}}$; (Mid) $\mathcal{C}_{\mathrm{inter}}$; (Right) $\mathcal{C}_{\mathrm{intra}}+\mathcal{C}_{\mathrm{inter}}$. We train Qwen2.5‑0.5B‑Instruct and LLaMA3.2-1b-Instruct on GSM8K using GRPO for 250 epochs. To accelerate observation, we use a training batch size of 16, generate 4 responses per question, and set the learning rate to 1e‑5. After each epoch, we perform pre-filling on the model on the entire dataset and record the angle concentration from the hidden states of the final layer output.
@@
A2. The angles between token hidden states show both intra-segment and inter-segment concentrations during inference. In particular, we should pay attention to the final layer as the inter‑segment clustering is most pronounced, resulting in the highest overall concentration.
@@ 2.3 How Can This Signal Be Leveraged to Accelerate Training?
Data‑wise Observation. Furthermore, to examine data-wise angle behavior during training, we track the model’s responses to samples with varying angle concentrations over epochs. As depicted in Fig. 5, surprisingly, the results revealed that during training, the model tends to prioritize learning from higher-angle concentration data before addressing lower-angle concentration data, which we introduce as Data‑wise Angle Concentration Pattern. For instance, by epoch 100, questions with maximal angular measurements were almost entirely answered correctly, whereas those with smaller angles remained uncorrected. Note that the angle concentration distribution is measured on the untrained model, indicating that despite its evolution during training, the initial angle concentration of the data provides meaningful guidance for the training process.
@@
Gradient-based Explanation. As shown in Eq. 4, gradients are influenced by both angle concentration and loss. Early in training, when losses are relatively uniform across samples, those with higher angle concentration receive stronger gradients and are learned faster. As training continues, angle concentration rises overall: high-angle samples, already mastered, see lower losses; while low‑angle samples gain concentration, inherit larger gradients, and are learned next. This creates a natural, angle-driven learning progression. Unlike traditional curriculum learning based on task difficulty, angle concentration offers a more intuitive, model-centric training signal.
@@ 3 GAIN-RL Framework
Probability Update Based on Accuracy and Angular Signals. The Gaussian mean $\mu_{t}$ sets the peak‑sampling region. We initialize $\mu_{0}=0$ to prioritize high‑angle concentration data, then gradually increase $\mu_{t}$ as the model masters these samples, shifting focus to lower‑angle concentration ones. At each step, $\mu_{t}$ is updated from the batch $d^{(t)}$ using its average accuracy and angle concentration,
@@
[2pt/2pt] Model Method
GSM8K
[3]
Math
[4]
AMC 23
[18]
AIME 24
[19]
Olympiad
Bench[20]
Minerva
Math[21]
Avg
Epo@
Same Acc
Speed
Up
Qwen 2.5
Math 1.5B
Instruct
@@
LLaMA 3.2
3B Instruct
@@
[…]
Qwen 2.5
Math 7B
Instruct
[…]
@@
[2pt/2pt] Model Method LCB [6] LCB Codeforces [22] Codeforces Humaneval+ [23] Avg Epo@ Speed
[2pt/2pt] Pass@1 Pass@8 Pass@1 Pass@8 Pass@1 Same Acc Up
Qwen 2.5 Coder
3B Instruct
@@
Appendix F experiments reveal that weighting the signal as $\mathcal{C}_{\mathrm{intra}}^{M}+k\cdot\mathcal{C}_{\mathrm{inter}}^{M}$ can further boost performance. For simplicity and broad applicability, we adopt the unweighted signal here and leave signal optimization to future work. Our main goal is to highlight angle concentration as a key signal.
Figure 7: Learning Dynamics of Different Methods on (Left) Qwen2.5‑Math‑1.5b‑Instruct and (Right) LLaMA‑3.2‑3b‑Instruct. The y‑axis shows average accuracy over GSM8K, Math, AMC 23, AIME 24, OlympiadBench, and Minerva Math. Performance is evaluated every 5 epochs.
@@
Figure 9: Fine‑tuning Performance on Single Tasks. (Left) GSM8k. (Mid) Math. (Right) AMC 23. Models are trained on the training sets and evaluated on their validation sets. ADARFT is excluded from GSM8K due to missing difficulty coefficients.
@@ 4.1 Training Efficiency of GAIN-RL
[…]
Figure 10: (Left) Ablation Study of Module Components. (Right) Small Batch Scalability Test. Experiments use Qwen2.5‑0.5b‑Instruct on Math training set, evaluate on the test set every 20 epochs.
@@ 4.2 Data Efficiency of GAIN-RL
Table 3: Performance of GAIN‑RL combined with PPO. Qwen2.5‑0.5b‑Instruct is trained and evaluated on three datasets. Implementation details are in Appendix E.
@@ 4.3 RL Algorithms Generalization
[…]
@@ 4.4 Performance on Individual Tasks
[…]
@@ 4.5 Ablation Studies
Small‑Batch Scalability Test. To validate the scalability of GAIN-RL under reduced training batch sizes, we evaluated the model’s performance with varying batch sizes. Specifically, we conducted experiments using the Qwen2.5-0.5b-instruct model on the Math dataset with training batch sizes of 512, 768, and 1024, respectively. The experimental outcomes, depicted in Fig. 10 (right), illustrate consistent and stable performance improvement across epochs for all batch sizes. Remarkably, even with the batch size reduced to half of the original size (n=512), which significantly reduces computational time and memory usage per epoch, GAIN-RL maintained performance comparable to vanilla GRPO. These results demonstrate that GAIN-RL effectively scales with smaller batch sizes, offering the flexibility to accelerate training by lowering batch size, with only minor performance degradation, especially beneficial under limited computational or memory resources.
@@ 5 Conclusion
[…]
@@ Appendix A Related Work
In parallel, data-centric strategies [28, 29, 30, 31, 32, 33] have emerged as promising alternatives for efficient fine-tuning. For example, [28] provides the first quantitative audit of preference datasets, introducing metrics for scale, noise, and information density that expose quality bottlenecks before any policy update. Direct Preference Optimization (DPO) [34] simplifies the entire loop by replacing on‑policy RL with a closed‑form classification loss, making the choice of high‑value preference data the primary driver of alignment quality. To further reduce annotation cost, Active Preference Optimization [35] casts RLHF as an active‑learning bandit, adaptively querying only the prompts expected to maximize reward‑model improvement. The latest data-centric strategies can be categorized based on their approach to data manipulation: data selection and data sequencing. Data selection techniques involve filtering extensive datasets to retain only a small subset of high-quality training data based on predefined metrics. Methods such as LIMO[12] and s1[13] have demonstrated that carefully curated small supervised fine-tuning datasets can achieve robust performance using orders of magnitude less data. On the other hand, data sequencing strategies[36] enhance model learning speed by rearranging the order of training data within existing datasets. Approaches like ADARFT[14] have shown that dynamically selecting data for each epoch can effectively accelerate the training process.
@@ B.1 Theoretical Justification of Angle-Dependent Effects in LLMs
Figure 11: Visualization of $W_{D}@W_{D}.T$ at different layers in Qwen2.5-0.5B-Instruct.
Figure 12: Visualization of $W_{O}@W_{O}.T$ at different layers in Qwen2.5-0.5B-Instruct.
Figure 13: Visualization of $W_{U}@W_{U}.T$ at different layers in Qwen2.5-0.5B-Instruct.
@@
For Observation 1, we show $W_{Q}@W_{K}.T$, $W_{V}@W_{V}.T$, $W_{D}@W_{D}.T$ of all different layers of Qwen2.5-0.5B-Instruct in Fig. 11, 12, and 13. It can be seen that different layers have this orthogonal relationship. In fact, the orthogonal relationship of matrices in neural networks has been studied since a long time ago. In particular, [37] proposed a new regularization method that encourages the weight matrix of the neural network to maintain orthogonality during training by introducing a self-orthogonality module. This method helps to improve the training stability and generalization ability of the model. [38] explores adding orthogonal regularization to weights during training to improve training stability. The author proposed an orthogonal regularization method for weights, aiming to solve the gradient vanishing and explosion problems encountered by deep convolutional neural networks during training. It can be seen that modules with orthogonality are found in various different models to improve the training stability and performance of the model. To the best of our knowledge, we are the first work to intuitively show this orthogonal performance in LLM, which can be more fully explored in subsequent research.
@@
Figure 14: Visualization of $cos(\angle(A_{i},A_{M}))$ and co-act neurons number at different layers in Qwen2.5-0.5B-Instruct.
@@
Justification. For a single token, suppose its input to the Attention block is $y$. Consider a sequence of tokens $y_{1},y_{2},\ldots,y_{M}]$, for token $y_{m}$ , its computation in the Attention layer can be expressed as:
@@
[…]